# Gravitational radiation from hyperbolic encounters in the presence of dark matter

### v1

@@ 6 Conclusion
In this work, motivated by its astrophysical significance and the prospect of its detection by future GW detectors and PTA [60], we studied the dynamics of hyperbolic encounters in the presence of the dark matter medium. Our approach to studying this involved computing radiation flux and the orbit dynamics of these encounters by treating the effect of the medium as a perturbative term in the Keplerian equations of motion. We observe that the power radiation due to the GW dominates over the radiation due to the dynamical friction generated by the dark matter halo. Next, while computing the change of eccentricity by computing various osculating elements, we observe again the effect of dynamical friction force felt by the secondary while interacting with the dark matter halo, and the effect of accretion force is negligible. But, the effect of the gravitational potential generated by the dark matter halo and GW backreaction effect is significantly more. There’s a genuine fall-off for eccentricity for the binaries due to these two effects for a certain time, while there is an eventual increase of the eccentricity due to the GW backreaction effect at late times. Interestingly, we observe that the change in eccentricity is more due to the dark matter potential for larger values of impact parameters. On the other hand, change in the eccentricity due to the GW backreaction increases with the decrease of impact parameter for a given value initial eccentricity. This can be attributed to the fact that, as we increase the initial impact parameter the secondary object gets to interact with the dark matter halo more, thereby increasing the effect of the dark matter potential in the change of the eccentricity. On the other hand, as we decrease the impact parameter, the GW backreaction effect is larger. These results are encouraging as the effect of the dark matter potential, which also depends on the parameter of the underlying model, provides a change in eccentricity and consequently can be constrained by observing such flyby events in future detectors.
@@
However, this analysis paves the way for only so much that can be done in these contexts. In doing the above study, we have made a number of simplifications, the foremost being the PN order in which it is to be computed. We have neglected the higher-order Post-Newtonian effects for these systems. For the flux calculation, we have only used the leading order PN results. The shape of the halo, which is taken here to be spherically symmetric, also needs to be relaxed, but we leave that for future work. Furthermore, while computing the osculating elements, we have neglected the evolution of the dark matter density profile. We have used only the static profile. But in general, the dark matter density profile evolves over time which is captured through the phase space distribution function [96, 97]. We like to include this in our future analysis as this will help us to make contact with a more realistic scenario. Furthermore, generalizing the analysis to include spinning binaries in this halo is also an interesting avenue to pursue in the future. Also, throughout our study, we have only focused on NFW dark matter profile. Recently, there have been other interesting dark matter profiles [135]. It will be interesting to extend our study to those cases. Last but not least, it will also be interesting to use the tools from the effective field theory to study this problem, perhaps along the line of [86, 136].
@@
Our analysis of osculating elements provides us with a platform to construct the GW waveform [60]. By solving the osculating equation, we can solve for the evolution of the binary’s orbital parameters, which in turn can be used to construct a waveform [105]. Then it will give us a route to understand and probe into the effects of dark matter contributions in such systems and possibly puts some constraints on the parameters using the GW data. In fact, following the analysis of [96], one can possibly comment on the dark matter environment by analyzing the dephasing of such GW waveform. We are currently in the process of constructing such a waveform for flyby events in the presence of dark matter and hope to report on that issue in the near future.
@@ Acknowledgements
Research of A.C. is supported by the Prime Minister’s Research Fellowship (PMRF-192002-1174) of the Government of India. A.B is supported by Mathematical Research Impact Centric Support Grant (MTR/2021/000490) by the Department of Science and Technology Science and Engineering Research Board (India), and Relevant Research Project grant (202011BRE03RP06633-BRNS) by the Board Of Research In Nuclear Sciences (BRNS), Department of Atomic Energy, India. We also thank the participants of the (virtual) workshop “Testing Aspects of General Relativity-II” and “New insights into particle physics from quantum information and gravitational waves” at Lethbridge University, Canada funded by McDonald Research Partnership-Building Workshop grant by McDonald Institute, where some parts of the results are presented by A.C., for valuable discussions.

### v2

@@ 7 Conclusion
In this work, motivated by its astrophysical significance and the prospect of its detection by future GW detectors and PTA [61], we studied the dynamics of hyperbolic encounters in the presence of the dark matter medium. Our approach to studying this involved computing radiation flux and the orbit dynamics of these encounters by treating the effect of the medium as a perturbative term in the Keplerian equations of motion. We observe that the power radiation due to the GW dominates over the radiation due to the dynamical friction generated by the dark matter halo. Next, while computing the change of eccentricity by computing various osculating elements, we observe again the effect of dynamical friction force felt by the secondary while interacting with the dark matter halo, and the effect of accretion force is negligible. But, the effect of the gravitational potential generated by the dark matter halo and GW backreaction effect is significantly more. There’s a genuine fall-off for eccentricity for the binaries due to these two effects for a certain time, while there is an eventual increase of the eccentricity due to the GW backreaction effect at late times. Interestingly, we observe that the change in eccentricity is more due to the dark matter potential for larger values of impact parameters. On the other hand, change in the eccentricity due to the GW backreaction increases with the decrease of impact parameter for a given value initial eccentricity. This can be attributed to the fact that, as we increase the initial impact parameter the secondary object gets to interact with the dark matter halo more, thereby increasing the effect of the dark matter potential in the change of the eccentricity. On the other hand, as we decrease the impact parameter, the GW backreaction effect is larger. These results are encouraging as the effect of the dark matter potential, which also depends on the parameter of the underlying model, provides a change in eccentricity and consequently can be constrained by observing such flyby events in future detectors. Furthermore, we extended our analysis by considering the effects of baryonic matter that may be present inside the halo. As discussed, it leads to the change of the profile, and we have considered one such model and analyzed the evolution of eccentricity. Curiously, it shows a sharper fall-off due to the extra damping provided by the presence of such matter.
@@
However, this analysis paves the way for only so much that can be done in these contexts. In doing the above study, we have made a number of simplifications, the foremost being the PN order in which it is to be computed. We have neglected the higher-order Post-Newtonian effects for these systems. For the flux calculation, we have only used the leading order PN results. The shape of the halo, which is taken here to be spherically symmetric, also needs to be relaxed, but we leave that for future work. Furthermore, while computing the osculating elements, we have neglected the evolution of the dark matter density profile. We have used only the static profile. But in general, the dark matter density profile evolves over time which is captured through the phase space distribution function [164, 163]. Also, the dark matter profile mentioned may not remain same during the entire duration of binary motion and merger [170]. It will be interesting to find a way to setup our computation done in this paper and investigate the evolution of eccentricity for that case. We like to include this in our future analysis as this will help us to make contact with a more realistic scenario. Furthermore, generalizing the analysis to include spinning binaries in this halo is also an interesting avenue to pursue in the future. Also, throughout our study, we have only focused on NFW dark matter profile. Recently, there have been other interesting dark matter profiles [172]. It will be interesting to extend our study to those cases. Last but not least, it will also be interesting to use the tools from the effective field theory to study this problem, perhaps along the line of [87, 173].
We have also looked for observational tools like braking index. The braking index is a useful tool for binaries at large separations and also in regimes where the environmental effects are dominant. We find that, it starts from a constant value when the bianries are in vacuum and start to decrease when the come close together inside the dark matter halo region. The study of eccentricity evolution complements it while also helping us to infer about environmental effects. Of course, our dissipative model was a simplistic one, $F=F_{0}r^{\gamma}v^{\delta}$, with more general forces having a dependence on spin, tidal deformability, etc. If we want to include more complex effects like halo feedbacks [163, 164], there can be a possibility of a new equilibrium[164] from the phase parameterization they develop. Eventually, one should go for studying relativistic and post-Newtonian effects in this framework, and this is something we would definitely like to take over in some future study.
@@
Our analysis of osculating elements provides us with a platform to construct the GW waveform [61]. By solving the osculating equation, we can solve for the evolution of the binary’s orbital parameters, which in turn can be used to construct a waveform [115]. Then it will give us a route to understand and probe into the effects of dark matter contributions in such systems and possibly puts some constraints on the parameters using the GW data. In fact, following the analysis of [164], one can possibly comment on the dark matter environment by analyzing the dephasing of such GW waveform. We are currently in the process of constructing such a waveform for flyby events in the presence of dark matter and hope to report on that issue in the near future. This will also help us to put constraints on the dark matter using multi-messenger astronomy similar to what has been done for the circular orbit [171]. Note that the analysis presented in this paper is valid for any mass ratios of the binary. It will be interesting to analyze the dephasing of gravitational waveform in the extreme mass-ratio limit [174]. This will be an interesting study in the context of gravitational wave phenomenology using future detectors like LISA.
@@ Acknowledgements
Research of A.C. is supported by the Prime Minister’s Research Fellowship (PMRF-192002-1174) of the Government of India. A.B AB is supported by the Core Research Grant (CRG/2023/005112) by the Department of Science and Technology Science and Engineering Research Board (India). R.K.S is supported by the Sabarmati Bridge Fellowship (MIS/IITGN-SBF/PHY/AB/2023-24/023) provided by Indian Institute of Technology Gandhinagar. We also thank the participants of the (virtual) workshop “Testing Aspects of General Relativity-II” and “New insights into particle physics from quantum information and gravitational waves” at Lethbridge University, Canada funded by McDonald Research Partnership-Building Workshop grant by McDonald Institute, where some parts of the results are presented by A.C., for valuable discussions. A.B also like to thank the Department of Physics and Astronomy of University of Lethbridge, especially Saurya Das and FISPAC Research Group, Department of Physics, University of Murcia, especially, Jose J. Fernández-Melgarejo for hospitality during the course of this work. A.B acknowledge associateship program of Indian Academy of Science, Bengaluru.